# SciAssess: Benchmarking LLM Proficiency in Scientific Literature Analysis

### v3

@@ 2.2.3 Biomedicine
Biomedical literature encompasses a wealth of specialized terminology and complex concepts, as well as a significant amount of non-textual information such as tables and figures. Effectively extracting and integrating these elements presents a crucial challenge. Given that tasks in the biomedical domain typically require precise identification and understanding of intricate biological entities, processes, and relationships, we have selected a set of representative tasks, including the recognition of specialized terminology, the comprehension of chart information, and the extraction of entity relationships, to evaluate the performance of LLMs in the biomedical field.
Biology ChartQA The analysis and understanding of biological properties, compositions, and processing techniques are critical for the discovery and development in life sciences. Often, this information is presented in charts, making it essential to extract and integrate such information with textual data. To assess the retrieval capabilities of models in the context of biological chart information, we have designed multiple-choice questions.
Chemical Entities Recognition This task involves recognizing chemical entity names using data from B5CDR [17] and additional expert-annotated data. It evaluates the performance of LLMs in identifying complex drug names.
Disease Entities Recognition Similarly, this task involves recognizing disease entity names using data from [17] and additional expert-annotated data, evaluating the performance of LLMs in identifying specialized disease names.
Compound Disease Recognition Proposed in B5CDR [17], this task evaluates the capability of LLMs to identify and understand associations between compounds and diseases.
Gene Disease Function & Gene Disease Regulation The Gene Disease Text Mining task focuses on "Gene-Disease" association semantics text mining. It evaluates the ability of models to extract and understand relationships between genes and diseases from scientific literature, with a focus on identifying gene and disease entities [18].
2.2.4 Drug Discovery
Drug discovery focuses on developing new therapeutics. Leveraging advanced intelligent tools, especially LLMs, can significantly enhance the efficiency and effectiveness of discovering and developing new drugs. These models are adept at parsing and analyzing complex data from diverse sources, including scientific texts and databases.
In this section, we examine LLMs in drug discovery, setting out a series of tasks designed to evaluate their ability to extract and process key information from literature related to pharmacology and molecular biology. These tasks aim to demonstrate how LLMs can be instrumental in identifying and understanding drug properties, interactions, and targets.
Affinity Extraction This task evaluates the LLM’s ability to extract an affinity table containing molecules’ tags, SMILES, and their affinities to different targets in bioassays. It tests several key capabilities of LLMs, including understanding complex and domain-specific language, as well as molecules and tables. Affinity data extraction requires not just surface-level text processing but also a deeper analysis to match different modalities.
Drug ChartQA The analysis of drug properties, compositions, and processing techniques is critical for drug discovery and development. Often, this information is presented in charts, making it essential to extract and integrate such information with textual data. To further assess the retrieval capabilities of models in the context of drug chart information, we have designed multiple-choice questions focusing on drug composition, processing methods, and properties.
Tag to Molecule This task evaluates the model’s ability to find the correct SMILES given its tag in a document. Typically, a molecule is shown with an chart of its structure and a tag below it. The LLM should recognize both the structure and the tag and understand their connection.
Markush to Molecule This task evaluates the model’s ability to obtain the correct SMILES given a Markush formula (in CXSMILES pattern) and its substituents.
Molecule in Document This task evaluates the model’s ability to determine whether a molecule (represented by SMILES) is mentioned in a document. The LLM should recognize all Markush formulas and their substituents, and then judge whether the required molecule is covered.
Reaction QA Organic and bio-catalyzed synthetic reactions are vital for the manufacture of drug-like molecules. Therefore, we designed a complex task to test the model’s capability in extracting information from schematic diagrams and texts of chemical reactions. The model is required to understand the charts specified in the articles and select the correct answer from the provided multiple-choice descriptions.
Drug Target Extraction This task evaluates the model’s ability to find all related targets (proteins or cell lines) of the bioassays given a document.
2.2.5 Organic Materials
Organic materials, derived from carbon-based molecules or polymers, exhibit distinct functionalities advantageous across a spectrum of applications. Distinguished from their inorganic counterparts by their modifiable properties and adaptability, these materials are vital in applications such as electronics, photonics, sensing, and energy. They leverage the vast potential of organic chemistry for technological advancements. In this section, our primary focus lies on two subfields within organic functional materials: organic electrolytes and polymer materials.
Electrolyte Table QA The composition and properties of organic electrolytes are crucial for battery performance, stability, and safety. To evaluate the model’s retrieval capabilities regarding electrolyte information, we posed multiple-choice questions about the components of solution systems and the dissolution reactions, focusing on their physical and chemical properties as presented in the tables within the articles.
OLED Property Extraction This task evaluates the LLM’s ability to extract information about OLED molecules and their optical properties. It tests several key capabilities, including their understanding of complex and domain-specific language and their ability to interpret and extract data from tables.
Polymer Chart QA The processing steps and properties of polymer materials are often represented through charts. Extracting information from these charts and integrating it with textual data is crucial. To further assess the retrieval capabilities of models concerning polymer chart information, we designed multiple-choice questions involving polymer composition, processing techniques, and properties.
Polymer Composition Extraction This task involves extracting the blend ratio of donor to acceptor in the most efficient solar cell from the text of scientific literature.
Polymer Property Extraction This task focuses on extracting vital values such as power conversion efficiency (PCE) and open-circuit voltage ($V_{\mathrm{OC}}$) from tables within the literature.
Solubility Extraction Organic electrolytes, extensively used in battery technologies, comprise organic solvents, lithium salts, and additives. Understanding solubility in organic electrolytes is crucial as it impacts the efficiency of electrolytic processes, product selectivity, and equipment design. This task evaluates the LLM’s capability in retrieving solubility-related tables. Papers typically select data from various aspects to describe the system, making it challenging to combine multiple tables for fuzzy matching. Therefore, we focus on examining the LLM’s semantic understanding ability, enabling the model to select the most relevant and comprehensive table related to “solubility” from numerous alternatives and convert it into the specified format.
Reaction Mechanism QA Investigating electrolyte reactions helps improve the solid electrolyte interphase (SEI) layer, which directly affects battery performance and lifespan. Studies in this area lead to the development of advanced electrolytes that enhance a robust SEI, resulting in more efficient and durable batteries. We design a complex task to test the capability of extracting information from schematic diagrams of chemical reaction mechanisms. LLM is required to understand the specified reaction diagram and select the correct answer from the provided multiple choices.
2.3 Data Quality, Privacy, and Copyright Compliance
@@
  • Expert Validation: Each original data point (as indicated by black tasks in Table 1) is independently labeled by two annotators who are domain experts in the relevant fields. If their labels agree, the label is accepted; if not, they engage in a discussion to determine the final label.
@@
  • Copyright Compliance: Our benchmark includes two types of data: some are adopted from existing benchmarks, and others are constructed from scratch by our team. For the data adopted from existing benchmarks, we provide the corresponding sources. For the data we created, we have obtained the necessary copyrights for the files used. To ensure full compliance with copyright laws, our repository only provides the Digital Object Identifier (DOI) for papers or patent number, and does not distribute the actual documents.
@@ 3.1 Experiment Setup
Baseline LLMs. To measure how state-of-the-art LLMs perform on SciAssess, we benchmark extensively:
Briefs about all models are provided in Appendix D.1.
Metrics. We choose the following metrics for each question type:
  • True/False & Multiple Choice: accuracy.
  • Table Extraction question: table recall, which is the recall rate for each value in the table.
  • Text Extraction question: recall.
  • Molecule Generation question: molecule similarity, which is measured by the Tanimoto similarity between two molecules’ Morgan fingerprints.
Experiment Workflow. For closed-source models, we utilize the official API calls provided by the model developers, while for open-source models, we obtain these models from HuggingFace [21], deploy them locally, and then perform the tests. For tasks requiring an article as context, the PDF content is converted to text and input into the LLM. If the LLM includes a built-in PDF parsing interface (e.g., Gemini and Moonshot), we simply use the interface; otherwise, we employ PyPDF2  [^8] https://pypdf2.readthedocs.io/en/3.x/, a widely-used open-source PDF parsing tool. Due to the input length limitations of the LLMs, tasks requiring the context of a PDF document are executed in a zero-shot manner. Tasks that do not require such context are evaluated using both zero-shot and few-shot methods.
@@ 3.2 Results and Analysis
In this section, we analyze the performance of LLMs across various scientific domains. The overall performance comparison of LLMs, as summarized in Table 2, reveals the distinct strengths and weaknesses of each model in science literature analysis.
Table 2: Performance Comparison of LLMs Across Various Scientific Domains. Orange and green indicate the best in closed and open source LLMs, respectively.
Domain Task ICL GPT-4o GPT-4 GPT-3.5 Moonshot Claude3 Doubao Gemini Llama3 DeepSeek Qwen2 Command $\text{R}^{\text{+}}$
Fundamental Science MMLU (science) 0-shot 0.839 0.783 0.629 0.774 0.795 0.720 0.799 0.766 0.737 0.782 0.647
3-shot 0.846 0.769 0.614 0.774 0.771 0.712 0.790 0.757 0.738 0.789 0.643
CMMLU (science) 0-shot 0.785 0.644 0.438 0.723 0.643 0.841 0.731 0.651 0.769 0.870 0.448
3-shot 0.785 0.646 0.432 0.728 0.631 0.833 0.736 0.658 0.768 0.867 0.455
Xiezhi-Ch (science) 0-shot 0.736 0.724 0.696 0.734 0.731 0.720 0.716 0.731 0.748 0.746 0.683
3-shot 0.736 0.708 0.690 0.732 0.706 0.706 0.723 0.736 0.726 0.745 0.672
Xiezhi-En (science) 0-shot 0.701 0.683 0.644 0.677 0.673 0.667 0.652 0.687 0.685 0.692 0.634
3-shot 0.699 0.670 0.641 0.679 0.658 0.650 0.654 0.683 0.665 0.697 0.632
Alloy Materials Alloy Chart QA 0-shot 0.533 0.600 0.333 0.333 0.400 0.467 0.667 0.467 0.333 0.400 0.200
Composition Extraction 0-shot 0.484 0.458 0.112 0.127 0.495 0.304 0.239 0.212 0.389 0.423 0.128
Temperature Extraction 0-shot 0.884 0.855 0.729 0.889 0.865 0.700 0.841 0.604 0.754 0.797 0.546
Sample Differentiation 0-shot 0.511 0.591 0.169 0.679 0.586 0.316 0.658 0.376 0.616 0.557 0.228
Treatment Sequence 0-shot 0.745 0.725 0.461 0.755 0.745 0.745 0.696 0.539 0.686 0.657 0.588
Biomedicine Biology Chart QA 0-shot 0.580 0.480 0.390 0.545 0.505 0.480 0.616 0.520 0.545 0.515 0.535
0-shot 0.454 0.665 0.540 0.201 0.844 0.911 0.678 0.400 0.536 0.832 0.850
Chemical Entities Recognition 3-shot 0.916 0.898 0.912 0.912 0.898 0.900 0.858 0.855 0.911 0.905 0.871
0-shot 0.279 0.765 0.153 0.000 0.653 0.675 0.437 0.526 0.331 0.722 0.258
Disease Entities Recognition 3-shot 0.822 0.849 0.879 0.785 0.782 0.811 0.807 0.787 0.825 0.826 0.647
0-shot 0.755 0.786 0.733 0.770 0.788 0.771 0.733 0.794 0.757 0.794 0.764
Compound Disease Recognition 3-shot 0.743 0.750 0.715 0.773 0.763 0.719 0.719 0.785 0.716 0.753 0.715
0-shot 0.931 0.974 0.864 0.771 0.944 0.779 0.954 0.996 0.819 0.930 0.884
Gene Disease Function 3-shot 0.945 0.927 0.896 0.845 0.931 0.772 0.868 0.876 0.830 0.814 0.888
0-shot 0.949 0.914 0.832 0.944 0.939 0.910 0.856 0.971 0.952 0.963 0.936
Gene Disease Regulation 3-shot 0.939 0.926 0.917 0.957 0.951 0.912 0.886 0.958 0.943 0.953 0.936
Drug Discovery Affinity Extraction 0-shot 0.072 0.042 0.025 0.040 0.097 0.050 0.040 0.064 0.017 0.075 0.043
Drug Chart QA 0-shot 0.333 0.400 0.067 0.400 0.200 0.533 0.533 0.400 0.400 0.400 0.533
Tag to Molecule 0-shot 0.040 0.022 0.000 0.016 0.035 0.094 0.169 0.034 0.014 0.000 0.031
Markush to Molecule 0-shot 0.634 0.632 0.429 0.462 0.644 0.217 0.218 0.478 0.543 0.358 0.332
3-shot 0.642 0.654 0.431 0.504 0.675 0.239 0.526 0.491 0.470 0.379 0.376
Molecule in Document 0-shot 0.580 0.700 0.500 0.460 0.480 0.560 0.640 0.680 0.460 0.460 0.460
Reaction QA 0-shot 0.705 0.674 0.442 0.253 0.663 0.442 0.305 0.611 0.368 0.442 0.316
Drug Target Identification 0-shot 0.721 0.791 0.526 0.607 0.794 0.622 0.768 0.600 0.687 0.410 0.485
Organic Materials Electrolyte Table QA 0-shot 0.940 0.790 0.370 0.670 0.870 0.710 0.880 0.460 0.720 0.620 0.450
OLED Property Extraction 0-shot 0.336 0.406 0.201 0.037 0.477 0.259 0.093 0.263 0.292 0.392 0.234
Polymer Chart QA 0-shot 0.800 0.667 0.400 0.800 0.467 0.867 0.800 0.867 0.733 0.933 0.800
Polymer Composition QA 0-shot 0.945 0.945 0.853 0.844 0.881 0.927 0.927 0.734 0.881 0.936 0.679
Polymer Property Extraction 0-shot 0.692 0.681 0.329 0.705 0.629 0.514 0.606 0.536 0.652 0.636 0.171
Solubility Extraction 0-shot 0.479 0.440 0.410 0.363 0.426 0.371 0.397 0.399 0.432 0.400 0.351
Reaction Mechanism QA 0-shot 0.545 0.636 0.455 0.545 0.455 0.636 0.727 0.500 0.545 0.591 0.591
@@ 3.2.1 Performances of Different Ability Levels
Table 3: Performances on Memorization (L1), Comprehension (L2), and Analysis & Reasoning (L3) tasks. Orange and green indicate the best in closed and open source LLMs, respectively.
Ability Level Question Type Metric GPT-4o GPT-4 GPT-3.5 Moonshot Claude3 Doubao Gemini Llama3 DeepSeek Qwen2 Command $\text{R}^{\text{+}}$
Memorization (L1) Multiple Choice Accuracy 0.77 0.70 0.59 0.73 0.69 0.73 0.73 0.71 0.72 0.77 0.60
Average Rank 1.88 6.50 10.50 4.50 7.38 7.12 5.25 4.88 5.75 1.75 10.50
Comprehension (L2) Molecule Generation Mol Similarity 0.04 0.02 0.00 0.02 0.04 0.09 0.17 0.03 0.01 0.00 0.03
Multiple Choice Accuracy 0.78 0.72 0.51 0.68 0.66 0.69 0.79 0.61 0.66 0.70 0.54
Table Extraction Table Accuracy 0.41 0.41 0.22 0.25 0.42 0.30 0.28 0.29 0.36 0.39 0.19
Text Extraction Recall 0.87 0.87 0.90 0.85 0.84 0.86 0.83 0.82 0.87 0.87 0.76
Average Rank 2.57 4.54 8.36 6.79 5.11 6.14 5.54 7.07 6.00 5.00 8.89
Analysis & Reasoning (L3) Molecule Generation Mol Similarity 0.64 0.65 0.43 0.50 0.68 0.24 0.53 0.49 0.47 0.38 0.38
Multiple Choice Accuracy 0.59 0.63 0.36 0.49 0.57 0.46 0.56 0.50 0.51 0.53 0.38
Text Extraction Recall 0.84 0.85 0.76 0.80 0.86 0.76 0.81 0.80 0.79 0.73 0.76
True/False Accuracy 0.66 0.71 0.48 0.61 0.61 0.65 0.67 0.61 0.57 0.56 0.52
Average Rank 4.20 3.45 8.50 5.35 3.95 7.10 5.45 5.50 6.95 7.10 8.45
@@
Table 3.2.1 presents the performance of evaluated LLMs across three progressive ability levels: Memorization (L1), Comprehension (L2), and Analysis & Reasoning (L3). Tasks are further categorized according to their question types, with average results and rankings provided for each ability levels. We observe the following: (1) Memorization (L1): GPT-4o and Qwen 2 demonstrates the highest average accuracy of 0.77, indicating consistently superior performance in memorization tasks. In contrast, GPT-3.5 shows the lowest performance with an accuracy of 0.59. (2) Comprehension (L2): GPT-4o excels in multiple-choice comprehension with an accuracy of 0.78 and maintains the top average rank of 2.57. Notably, the only molecule generation task at the L2 level, Tag to Molecule, reveals poor performance across all LLMs. As illustrated in Figure 4, current PDF parsing technologies, whether open-source like PyPDF or proprietary like Gemini or Moonshot, fail to effectively parse molecular structures in documents. Consequently, these technologies struggle with the Tag to Molecule task. We propose that a critical advancement for future LLM-based literature analysis tools is the integration of PDF parsing solutions capable of recognizing molecular structures. (3) Analysis & Reasoning (L3): The average rank reveals GPT-4, Claude3, GPT-4o and are the top performers with ranks of 3.45, 3.95, and 4.20, respectively.
Overall, GPT-4o consistently ranks high across all ability levels, particularly excelling in memorization tasks. GPT-4 and Claude3 also demonstrate strong overall performance, especially in comprehension and reasoning. GPT-3.5, despite being an earlier model, shows limitations compared to its successors. Command $\text{R}^{\text{+}}$, while competitive in some areas, generally underperforms compared to other models.
Based on these observations, we suggest the following recommendations: (1) For tasks heavily reliant on memorization, GPT-4o and Qwen2 are recommended due to their high accuracy and ranking; (2) For comprehension tasks, particularly those involving complex data extraction and generation, GPT-4o is ideal choices. (3) For analysis and reasoning tasks, Claude3 and GPT-4 provide reliable performance and should be considered.

### v5

@@ 2.2.1 Biology
Biological literature encompasses a wealth of specialized terminology and complex concepts, as well as a significant amount of non-textual information such as tables and figures. Effectively extracting and integrating these elements presents a crucial challenge. Given that tasks in the biological domain typically require precise identification and understanding of intricate biological entities, processes, and relationships, we have selected a set of representative tasks, including the recognition of specialized terminology, the comprehension of chart information, and the extraction of entity relationships, to evaluate the performance in this field.
In this domain, following tasks are devised: MMLU-Pro-Biology, biology chart QA, chemical entities recognition, compound disease recognition, disease entities recognition, and gene disease function. Detailed descriptions and prompts are provided in Appendix C.1.
2.2.2 Chemistry
The field of chemistry involves a vast array of complex molecular structures, chemical reactions, and properties, alongside a substantial amount of data presented in formulas, reaction equations, and diagrams. Effectively processing and interpreting these components is a significant challenge for language models. Tasks in the chemical domain demand precise understanding of molecular compositions, reaction mechanisms, and material properties. To evaluate the performance of LLMs in this domain, we have selected representative tasks such as the recognition of chemical compounds, the interpretation of reaction pathways, and the extraction of relationships between chemical entities.
We devise following tasks for organic materials: MMLU-Pro-Chemistry, electrolyte table QA, OLED property extraction, polymer chart QA, polymer composition extraction, polymer property extraction, solubility extraction, reactant QA, and reaction mechanism QA. Detailed descriptions and prompt templates are provided in Appendix C.2.
2.2.3 Materials
Materials science encompasses a broad range of substances, including metals, ceramics, polymers, and composites, each with distinct properties and applications. These materials are widely used across industries such as aerospace, automotive, and construction. By fine-tuning their composition, structure, and processing techniques, materials can be engineered to meet specific performance requirements Caron and Khan (1983). Accurately extracting material compositions, structural characteristics, and process parameters from the literature is essential for advancing material design and optimization.
Specifically, following tasks are devised: material QA, Alloy Chart QA, composition extraction, temperature QA, sample differentiation, and treatment sequence. Detailed descriptions and prompt templates are provided in Appendix C.3.
2.2.4 Medicine
Medicine focuses on developing new therapeutics. Leveraging advanced intelligent tools, especially LLMs, can significantly enhance the efficiency and effectiveness of discovering and developing new drugs. To evaluate the capability of LLMs in this domain, it is imperative to develop specialized tasks that reflect the complexities and nuances of biomedical research. By designing targeted tasks, we can better assess the ability of LLMs to navigate and interpret the wealth of information critical to the development of new therapeutics.
Specifically, we devise: MMLU-Pro-Health, affinity extraction, drug chart QA, tag to molecule, markush to molecule, molecule in document. Detailed descriptions and prompts are provided in Appendix C.4.
2.3 Data Quality, Privacy, and Copyright
@@
Distractor Construction: Our data points are human-annotated, as well as the distractors. And how the distractors are determined depends on the specific task. For example, in value-type multiple-choice questions, the distractors are values near the ground truth. For extraction tasks, the distractors are other targets in the given context, except for the ground truth target.
Expert Validation: Each data point (as indicated by black tasks in Table 1) is independently labeled by two annotators who are domain experts in the relevant fields. If their labels agree, the label is accepted; if not, they engage in a discussion to determine the final label. Their initial annotations have a Cohen’s Kappa value McHugh (2012) of 0.75, which indicates high reliability or agreement.
@@
Copyright Compliance: Our benchmark includes two types of data: some are adopted from existing benchmarks, and others are constructed from scratch by our team. For the data adopted from existing benchmarks, we provide the corresponding sources. For the data we created, we have obtained the necessary copyrights for the files used. To ensure full compliance with copyright laws, our repository only provides the Digital Object Identifier (DOI) for papers or patent number, and does not distribute the actual documents. Researchers need to download the necessary files independently. Detailed instructions is included in the codebase to guide researchers on how to place the downloaded documents into the designated folder.
@@ 3.1 Experiment Setup
Domain Task o1 GPT-4o GPT-4 GPT-3.5 Moonshot Claude3 Doubao Gemini Llama3.1 Qwen2.5 Mixtral
Biology MMLU-Pro-Biology* 0.901 0.874 0.845 0.650 0.755 0.781 0.770 0.842 0.815 0.840 0.743
Biology Chart QA 0.653 0.558 0.442 0.312 0.518 0.402 0.523 0.548 0.477 0.487 0.422
Chemical Entities Recognition* 0.862 0.795 0.817 0.649 0.821 0.815 0.749 0.745 0.836 0.764 0.707
Compound Disease Recognition* 0.745 0.733 0.753 0.636 0.745 0.737 0.733 0.751 0.768 0.712 0.757
Disease Entities Recognition* 0.831 0.763 0.670 0.688 0.654 0.684 0.742 0.767 0.793 0.793 0.737
Gene Disease Function* 0.687 0.410 0.587 0.391 0.538 0.506 0.539 0.558 0.474 0.438 0.418
Chemistry MMLU-Pro-Chemistry* 0.868 0.745 0.621 0.303 0.428 0.496 0.446 0.683 0.676 0.723 0.501
Electrolyte Table QA 0.925 0.855 0.810 0.305 0.765 0.435 0.745 0.765 0.755 0.785 0.455
OLED Property Extraction 0.394 0.438 0.455 0.280 0.160 0.055 0.413 0.419 0.563 0.499 0.355
Polymer Chart QA 1.000 0.867 0.733 0.667 0.800 0.467 0.400 0.733 0.800 0.800 0.800
Polymer Composition QA 0.986 0.938 0.947 0.330 0.971 0.555 0.804 0.947 0.852 0.914 0.493
Polymer Property Extraction 0.606 0.759 0.758 0.562 0.736 0.634 0.508 0.580 0.690 0.692 0.573
Solubility Extraction 0.427 0.444 0.431 0.408 0.445 0.375 0.409 0.440 0.447 0.437 0.314
Reactant QA 0.559 0.487 0.441 0.272 0.415 0.241 0.272 0.344 0.385 0.379 0.231
Reaction Mechanism QA 0.682 0.591 0.273 0.227 0.500 0.136 0.409 0.682 0.455 0.500 0.273
Material Material QA 0.821 0.768 0.722 0.521 0.620 0.620 0.669 0.722 0.738 0.719 0.631
Alloy Chart QA 0.533 0.467 0.200 0.400 0.333 0.467 0.533 0.600 0.467 0.533 0.467
Composition Extraction 0.488 0.462 0.467 0.189 0.423 0.427 0.398 0.389 0.457 0.430 0.177
Temperature QA 0.836 0.807 0.734 0.295 0.845 0.353 0.488 0.715 0.652 0.647 0.382
Sample Differentiation 0.392 0.624 0.595 0.329 0.688 0.245 0.376 0.586 0.624 0.578 0.426
Treatment Sequence 0.624 0.594 0.678 0.485 0.683 0.480 0.658 0.634 0.614 0.658 0.673
Medicine MMLU-Pro-Health* 0.784 0.763 0.715 0.531 0.644 0.614 0.605 0.663 0.710 0.685 0.603
Affinity Extraction 0.068 0.101 0.076 0.055 0.063 0.045 0.081 0.052 0.047 0.071 0.049
Drug Chart QA 0.600 0.467 0.333 0.333 0.333 0.467 0.400 0.533 0.400 0.333 0.400
Tag2Mol 0.127 0.229 0.092 0.023 0.133 0.061 0.105 0.211 0.143 0.136 0.021
Markush2Mol 0.662 0.585 0.684 0.523 0.391 0.503 0.565 0.683 0.425 0.443 0.576
Mol In Document 0.840 0.600 0.620 0.440 0.480 0.640 0.560 0.580 0.520 0.600 0.460
Table 2: Performance comparison of LLMs across various scientific domains. Orange and green indicate the best in closed and open source LLMs, respectively. Chain-of-thought prompt is implemented for each task and model, except for OpenAI-o1. * indicates 3-shot.
Baseline LLMs. To measure how leading LLMs perform on SciAssess, we benchmark extensively. For closed-source LLMs, we test OpenAI-o1, GPT-4o, GPT-4, GPT-3.5, Gemini-1.5-Pro, Claude 3 Opus, Moonshot-v1 and Doubao. For open-source LLMs, we test Llama-3.1-70B, Mixtral-8x22B-Instruct-v0.1, and Qwen-2.5-72B. Briefs about all models are provided in Appendix D.
Experiment Workflow. For closed-source models, we utilize the official API calls provided by the model developers, while for open-source models, we obtain these models from HuggingFace Wolf et al. (2019), deploy them locally with vllm Kwon et al. (2023), and then perform the tests. Tasks curated by us require real context from papers, thus the PDF content needs to be converted to text as inputs for LLMs. If the LLM includes a built-in PDF parsing interface (e.g., Gemini and Moonshot), we simply use the interface; otherwise, we employ PyPDF2  [^2] https://pypdf2.readthedocs.io/en/3.x/, a widely-used open-source PDF parsing tool. Our aim is to explore the ability boundary of LLMs, thus strategies that enhance LLMs’ inference ability (i.e., in-context learning Brown (2020) and chain-of-thought Wei et al. (2022)) are adopted. Specifically, due to the input length limitations of the LLMs, tasks requiring long context of a PDF document are executed in a zero-shot manner. Tasks that do not require such long context (e.g., MMLU-Pro, entities recognition) are evaluated using 3-shot settings. And chain-of-thought prompt is implemented in every task by prompting the model to think step-by-step before concluding. The only exception is OpenAI-o1, whose official prompt guideline suggests users to “avoid chain-of-thought prompts”. We also provide complete performance evaluation results with chain-of-thought prompts in Appendix E.
@@ 3.2 Results and Analysis
In this section, we analyze the performance of LLMs on SciAssess. The overall performance comparison, as summarized in Table 2, reveals the distinct strengths and weaknesses of each model in science literature analysis.
@@ 3.2.1 Performances of Different Ability Levels
Ability Level Question Type Metric o1 GPT-4o GPT-4 GPT-3.5 Moonshot Claude3 Doubao Gemini Llama3.1 Qwen2.5 Mixtral
Memorization (L1) Multiple Choice Accuracy 0.843 0.788 0.726 0.501 0.612 0.628 0.622 0.728 0.735 0.742 0.619
Average Rank 1.000 2.000 6.000 11.000 10.000 7.000 8.000 5.000 4.000 3.000 9.000
Comprehension (L2) Multiple Choice Accuracy 0.790 0.708 0.600 0.377 0.652 0.449 0.556 0.692 0.629 0.643 0.488
Table Extraction Recall 0.397 0.441 0.437 0.299 0.365 0.307 0.362 0.376 0.441 0.426 0.294
Text Extraction F1-score 0.781 0.675 0.707 0.591 0.690 0.686 0.691 0.705 0.718 0.677 0.655
Mol. Generation Mol. Similarity 0.127 0.229 0.092 0.023 0.133 0.061 0.105 0.211 0.143 0.136 0.021
Average Rank 3.250 3.375 5.250 10.500 5.500 8.750 7.000 3.750 3.125 5.250 10.250
Analysis & Reasoning (L3) Multiple Choice Accuracy 0.544 0.567 0.436 0.276 0.534 0.207 0.352 0.537 0.488 0.486 0.310
Mol. Generation Mol. Similarity 0.662 0.585 0.684 0.523 0.391 0.503 0.565 0.683 0.425 0.443 0.576
True/False Accuracy 0.732 0.597 0.649 0.462 0.582 0.560 0.609 0.607 0.567 0.629 0.566
Average Rank 2.000 3.667 3.333 9.333 7.333 9.667 6.000 3.333 7.667 6.000 7.667
Table 3: Performance on Memorization (L1), Comprehension (L2), and Analysis & Reasoning (L3) tasks.
@@
Table 3 presents the performance of evaluated LLMs across three progressive ability level. Tasks are further categorized according to their question types, with average results and rankings provided for each ability levels. We observe the following: (1) Memorization (L1): OpenAI-o1 and Qwen-2.5 demonstrates the highest average accuracy of 0.843 and 0.742, respectively, indicating consistently superior performance in memorization tasks. (2) Comprehension (L2): OpenAI-o1 excels in multiple-choice and text extraction comprehension with accuracy of 0.790 and 0.781, respectively, and maintains the top average rank of 3.25. Notably, the only L2-level molecule generation task, Tag to Molecule, reveals poor performance across all LLMs. As illustrated in Figure 4, current PDF parsing technologies, whether open-source like PyPDF or proprietary like Gemini or Moonshot, fail to effectively parse molecular structures in documents. Consequently, LLMs struggle with the Tag to Molecule task. We propose that a critical advancement for future LLM-based literature understanding assistant is the integration of PDF parsing solutions capable of recognizing molecular structures. (3) Analysis & Reasoning (L3): The average rank reveals OpenAI-o1, GPT-4, and Gemini are the top performers with ranks of 2.00, 3.33, and 3.33, respectively.
Overall, OpenAI-o1 consistently ranks high across all ability levels. GPT-4o and Gemini also demonstrate strong overall performance, especially in memorization and reasoning.
Based on these observations, we suggest the following recommendations: (1) For tasks heavily reliant on memorization, OpenAI-o1 and Qwen2.5 are recommended due to their high accuracy and ranking; (2) For comprehension tasks, particularly those involving complex data extraction and generation, OpenAI-o1 is an ideal choice. (3) For analysis and reasoning tasks, OpenAI flagship models (i.e., o1, GPT4, GPT-4o), and Gemini provide reliable performance and should be considered.